I messed up with the last question i submitted.
We'll say that \ax+b\ means the square root of that.
\5x+6\-2= -\3x+4\ solve for x. please help me with the steps to solving it

Answer by snwbrdfrk(10)

ok so the equation is \5x+6\-2=-\3x+4\
if you want to solve for x here is an easy thing to do.
i dont know if you have learned this or not but this is what i do when this happens and i know that it is aloud.
when you have an equation like that multiply everything by the squareroot of 2
so it should look like this
\5x+6\^2 -2^2=-\3x+4\^2
it should look like this now
5x+6 +4 = 3x+4
now we have to find x but we need to simplify a little.
5x + 10 = 3x + 4
now subtract both sides by 4 to get it to the other side
5x + 6 = 3x
now subtract both sides by 5x
6= 3x-5x
6= -2x
divide both sides by 2
6/-2=x
answer is -3